# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ A-A Proof of Theorem 1
$\leq\alpha_{\mathcal{E}_{0}}^{2}\cdot H.$ (47)
@@
$\lim_{t\to\infty}\mathbb{E}[\mathcal{T}^{t}]=\lim_{t\to\infty}\mathbb{E}[\mathbb{E}[\mathcal{T}^{t}|\mathcal{E}_{0}]]=\mathbb{E}[\lim_{t\to\infty}\mathbb{E}[\mathcal{T}^{t}|\mathcal{E}_{0}]]\leq\mathbb{E}[\alpha_{\mathcal{E}_{0}}^{2}H]=\alpha^{2}H.$ (48)
Step 5: By Eqn. (48), we derive that there exists $T_{0}>0$ such that $\mathbb{E}[\mathcal{T}^{t}]\leq 2\alpha^{2}H$ for $t>T_{0}$. Plugging into Eqn. (40) and then taking $\lambda=\sqrt{\alpha^{2}TH/{\log(K(\epsilon))}}$, we obtain
@@
$\leq\alpha\sqrt{TH\log(K(\epsilon))}+T\epsilon+T_{0},$ (49)
which finishes the proof of Theorem 1. ∎
A-B Proof of Theorem 2
Recall that
$V_{1,\pi}^{\mathcal{E}}\big{(}s_{1}^{\ell}\big{)}-V_{1,\pi}^{\mathcal{E}^{\prime}}\big{(}s_{1}^{\ell}\big{)}$ $=\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}\mathbb{E}_{s^{\prime}\sim P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}}\left[V_{h+1,\pi}^{\mathcal{E}}\left(s^{\prime}\right)\right]-\mathbb{E}_{s^{\prime}\sim P_{h}^{\mathcal{E}^{\prime}}\left(\cdot\mid s_{h},a_{h}\right)}\big{[}V_{h+1,\pi}^{\mathcal{E}}\big{(}s^{\prime}\big{)}\big{]}\bigg{]}$ (50)
@@ A-B Proof of Theorem 2
$=\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}-P_{h}^{\mathcal{E}^{\prime}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}\bigg{]}$
@@
$\leq\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}-P_{h}^{\mathcal{E}^{\prime}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}\bigg{]}$
$\quad+\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}\|V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}-V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}\|_{2}\bigg{]}+\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\left[r_{h}^{\mathcal{E}}\left(s_{h},a_{h}\right)-r_{h}^{\mathcal{E}^{\prime}}\left(s_{h},a_{h}\right)\right]$
Since $P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}\in[0,H]$, we can divide the value range $[0,H]$ evenly into $\frac{3H^{2}}{\epsilon}$ parts. For each $(s,a,h)$, we construct a covering set $\{\mathcal{I}_{sah}^{1},...,\mathcal{I}_{sah}^{m}\}$ for $[0,H]$ where $m=\frac{3H^{2}}{\epsilon}$. Each set is of length $\frac{\epsilon}{3H}$. Since $\|V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}\|_{2}\in[0,H\sqrt{S}]$, we construct a covering set $\{\mathcal{J}_{h}^{1},...,\mathcal{J}_{h}^{m^{\prime}}\}$ for $[0,H\sqrt{S}]$ where $m^{\prime}=\frac{6H^{2}\sqrt{S}}{\epsilon}$. For reward function, we divide the value range $[0,1]$ evenly into $\frac{2H}{\epsilon}$ parts for all $(s,a)\in\mathcal{S}\times\mathcal{A},h\in[H]$. The covering set is $\{\mathcal{C}_{1},...,\mathcal{C}_{n}\}$ where $n=\frac{3H}{\epsilon}$. Then, we construct the partition $\{\Theta_{k}\}_{k=1}^{K}$ that $\mathcal{E}\in\Theta_{k}$ if for any $s,a,h$,
@@
Since $K(\epsilon)\leq(\frac{3H^{2}}{\epsilon})^{SAH}\cdot(\frac{6H^{2}\sqrt{S}}{\epsilon})^{H}\cdot(\frac{3H}{\epsilon})^{SAH}$, we have
$\log(K(\epsilon))\leq SAH\log(\frac{3H^{2}}{\epsilon})+H\log(\frac{6H^{2}\sqrt{S}}{\epsilon})+SAH\log(\frac{3H}{\epsilon})\leq 3SAH\log(\frac{6H^{2}\sqrt{S}}{\epsilon}).$
From Theorem 1, we have
@@ A-C Proof of Theorem 3
Then we construct the partition of $\Theta$ as following: $\mathcal{E}$ and $\mathcal{E}^{\prime}$ belong to the same partition if and only if $(\psi_{h}^{P,\mathcal{E}})_{i}$ and $(\psi_{h}^{P,\mathcal{E}^{\prime}})_{i}$ belong to the same partition of $\mathcal{F}$, $\forall i\in[d]$.
Recall that
@@
$\leq H\sum_{h=1}^{H}\ell_{1}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{E}^{\prime}})+\sum_{h=1}^{H}\ell_{1}(R_{h}^{\mathcal{E}},R_{h}^{\mathcal{E}^{\prime}}).$
Next, we use the coverage of $\mathcal{F}$ under $\ell_{g}$ to bound $\ell_{1}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{E}^{\prime}})$ and $\ell_{1}(R_{h}^{\mathcal{E}},R_{h}^{\mathcal{E}^{\prime}})$.
$\parallel\phi_{h}^{P}(s,a)\parallel_{2}\leq 1\Rightarrow|\phi_{h}^{P}(s,a)_{i}|\leq 1,\forall i\in[d]$. For any $\mathcal{E},\mathcal{E}^{\prime}$ that belong to the same partition, we have
$\ell_{1}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{E}^{\prime}})$ $=\sup_{s,a}\int_{s^{\prime}}|P_{h}^{\mathcal{E}}(s^{\prime}|s,a)-P_{h}^{\mathcal{E}^{\prime}}(s^{\prime}|s,a)|$ (52)
@@ A-D Proof of Proposition 1
$r^{\prime}_{h}(s,a)=r_{h}(s,a)+\frac{\lambda}{2}\mathbb{E}_{t}\left[D_{\mathrm{KL}}\left((P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\right].$ (53)
@@
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\cdot\epsilon\cdot\left(1+2\log\frac{1}{\beta}\right).$ (54)
@@
$\bigg{|}D_{\text{KL}}\left((P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)-D_{\text{KL}}\left((P_{h}^{\mathcal{E}}\otimes R_{h}^{\mathcal{E}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\bigg{|}$
$\leq\bigg{|}\int_{\mathcal{S}}P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x-\int_{\mathcal{S}}P_{h}^{\mathcal{E}}(x|s_{h},a_{h})\log\frac{P_{h}^{\mathcal{E}}(x|s_{h},a_{h})}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x\bigg{|}$
$+\bigg{|}\int_{[0,1]}R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x-\int_{[0,1]}R_{h}^{\mathcal{E}}(x|s_{h},a_{h})\log\frac{R_{h}^{\mathcal{E}}(x|s_{h},a_{h})}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x\bigg{|}.$ (55)
Let $o=(s_{h},a_{h})$. For the first term in Eqn. (A-D), we have the following bound
$\bigg{|}\int_{\mathcal{S}}P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\mathcal{E}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
$\leq\bigg{|}\int_{\mathcal{S}}P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}+\bigg{|}\int_{\mathcal{S}}P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\mathcal{E}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
@@
$\leq\int_{\mathcal{S}}\left|P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)-P_{h}^{\mathcal{E}}(x|o)\right|\cdot\bigg{|}\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\bigg{|}\mathrm{d}x+\int_{\mathcal{S}}B\cdot\bigg{|}\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
@@
$\leq B\cdot\left(1+2\log\frac{1}{\beta}\right)\int_{\mathcal{S}}\bigg{|}\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
$\leq\frac{(1-2\log{\beta})\epsilon}{2H^{2}},$ (56)

### v3

@@ Appendix A Proofs of Theorem and Proposition
$\leq\alpha_{\mathcal{E}_{0}}^{2}\cdot H.$ (A.22)
@@
$\lim_{t\to\infty}\mathbb{E}[\mathcal{T}^{t}]=\lim_{t\to\infty}\mathbb{E}[\mathbb{E}[\mathcal{T}^{t}|\mathcal{E}_{0}]]=\mathbb{E}[\lim_{t\to\infty}\mathbb{E}[\mathcal{T}^{t}|\mathcal{E}_{0}]]\leq\mathbb{E}[\alpha_{\mathcal{E}_{0}}^{2}H]=\alpha^{2}H.$ (A.23)
Step 5: By Eqn. (A.23), we derive that there exists $T_{0}>0$ such that $\mathbb{E}[\mathcal{T}^{t}]\leq 2\alpha^{2}H$ for $t>T_{0}$. Plugging into Eqn. (A.15) and then taking $\lambda=\sqrt{\alpha^{2}TH/{\log(K(\epsilon))}}$, we obtain
@@
$\leq\alpha\sqrt{TH\log(K(\epsilon))}+T\epsilon+T_{0},$ (A.24)
which finishes the proof of Theorem 4.11. ∎
A.2 Proof of Theorem 4.15
$V_{1,\pi}^{\mathcal{E}}\big{(}s_{1}^{\ell}\big{)}-V_{1,\pi}^{\mathcal{E}^{\prime}}\big{(}s_{1}^{\ell}\big{)}$ (A.25)
$\overset{(a)}{=}\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}\mathbb{E}_{s^{\prime}\sim P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}}\left[V_{h+1,\pi}^{\mathcal{E}}\left(s^{\prime}\right)\right]-\mathbb{E}_{s^{\prime}\sim P_{h}^{\mathcal{E}^{\prime}}\left(\cdot\mid s_{h},a_{h}\right)}\big{[}V_{h+1,\pi}^{\mathcal{E}}\big{(}s^{\prime}\big{)}\big{]}\bigg{]}$
@@
$\overset{(b)}{=}\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}-P_{h}^{\mathcal{E}^{\prime}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}\bigg{]}$
@@
$\overset{(c)}{\leq}\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}-P_{h}^{\mathcal{E}^{\prime}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}\bigg{]}$
$\quad+\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}\|V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}-V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}\|_{2}\bigg{]}+\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\left[r_{h}^{\mathcal{E}}\left(s_{h},a_{h}\right)-r_{h}^{\mathcal{E}^{\prime}}\left(s_{h},a_{h}\right)\right],$
where (a) uses Lemma C.1, (b) adds and subtracts the term $P_{h}^{\mathcal{E}^{\prime}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}^{\prime}}\big{(}\cdot\big{)}$, (c) uses the Cauchy-schwartz inequality.
Since $P_{h}^{\mathcal{E}}\big{(}\cdot\mid s_{h},a_{h}\big{)}^{T}V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}\in[0,H]$, we can divide the value range $[0,H]$ evenly into $\frac{3H^{2}}{\epsilon}$ parts. For each $(s,a,h)$, we construct a covering set $\{\mathcal{I}_{sah}^{1},...,\mathcal{I}_{sah}^{m}\}$ for $[0,H]$ where $m=\frac{3H^{2}}{\epsilon}$. Each set is of length $\frac{\epsilon}{3H}$. Since $\|V_{h+1,\pi}^{\mathcal{E}}\big{(}\cdot\big{)}\|_{2}\in[0,H\sqrt{S}]$, we construct a covering set $\{\mathcal{J}_{h}^{1},...,\mathcal{J}_{h}^{m^{\prime}}\}$ for $0,H\sqrt{S}]$ where $m^{\prime}=\frac{6H^{2}\sqrt{S}}{\epsilon}$. For reward function, we divide the value range $[0,1]$ evenly into $\frac{2H}{\epsilon}$ parts for all $(s,a)\in\mathcal{S}\times\mathcal{A},h\in[H]$. The covering set is $\{\mathcal{C}_{1},...,\mathcal{C}_{n}\}$ where $n=\frac{3H}{\epsilon}$. Then, we construct the partition $\{\Theta_{k}\}_{k=1}^{K}$ that $\mathcal{E}\in\Theta_{k}$ if for any $s,a,h$,
@@
Since
$K(\epsilon)\leq(\frac{3H^{2}}{\epsilon})^{SAH}\cdot(\frac{6H^{2}\sqrt{S}}{\epsilon})^{H}\cdot(\frac{3H}{\epsilon})^{SAH},$
we have
$\log(K(\epsilon))\leq SAH\log\bigg{(}\frac{3H^{2}}{\epsilon}\bigg{)}+H\log\bigg{(}\frac{6H^{2}\sqrt{S}}{\epsilon}\bigg{)}+SAH\log\bigg{(}\frac{3H}{\epsilon}\bigg{)}\leq 3SAH\log\bigg{(}\frac{6H^{2}\sqrt{S}}{\epsilon}\bigg{)}.$
From Theorem 4.11, we have
@@
Then we construct the partition of $\Theta$ as following: $\mathcal{E}$ and $\mathcal{E}^{\prime}$ belong to the same partition if and only if $(\psi_{h}^{P,\mathcal{E}})_{i}$ and $(\psi_{h}^{P,\mathcal{E}^{\prime}})_{i}$ belong to the same partition of $\mathcal{F}$, $\forall i\in[d]$. Then, we have
@@
$\overset{(a)}{=}\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}\bigg{[}\mathbb{E}_{s^{\prime}\sim P_{h}^{\mathcal{E}}(\cdot|s_{h},a_{h})}\left[V_{h+1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim P_{h}^{\mathcal{E}^{\prime}}(\cdot|s_{h},a_{h})}\left[V_{h+1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s^{\prime})\right]\bigg{]}+\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}\left[R_{h}^{\mathcal{E}}(s_{h},a_{h})-R_{h}^{\mathcal{E}^{\prime}}(s_{h},a_{h})\right]$
@@
$\overset{(c)}{\leq}H\sum_{h=1}^{H}\ell_{1}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{E}^{\prime}})+\sum_{h=1}^{H}\ell_{1}(R_{h}^{\mathcal{E}},R_{h}^{\mathcal{E}^{\prime}}),$ (A.26)
where (a) uses Lemma C.1, (b) follows from $V_{h+1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s^{\prime})\leq H$. Next, we use the coverage of $\mathcal{F}$ under $\ell_{g}$ to bound $\ell_{1}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{E}^{\prime}})$ and $\ell_{1}(R_{h}^{\mathcal{E}},R_{h}^{\mathcal{E}^{\prime}})$.
First notice that
$\parallel\phi_{h}^{P}(s,a)\parallel_{2}\leq 1\Rightarrow|\phi_{h}^{P}(s,a)_{i}|\leq 1,\forall i\in[d].$
For any $\mathcal{E},\mathcal{E}^{\prime}$ that belong to the same partition, we have
$\ell_{1}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{E}^{\prime}})$ $=\sup_{s,a}\int_{s^{\prime}}|P_{h}^{\mathcal{E}}(s^{\prime}|s,a)-P_{h}^{\mathcal{E}^{\prime}}(s^{\prime}|s,a)|$ (A.27)
@@
$r^{\prime}_{h}(s,a)=r_{h}(s,a)+\frac{\lambda}{2}\mathbb{E}_{t}\left[D_{\mathrm{KL}}\left((P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\right].$ (A.28)
@@
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\cdot\epsilon\cdot\left(1+\log\frac{B}{\beta}\right).$ (A.29)
@@
$\bigg{|}D_{\text{KL}}\left((P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)-D_{\text{KL}}\left((P_{h}^{\mathcal{E}}\otimes R_{h}^{\mathcal{E}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\bigg{|}$
$\leq\bigg{|}\int_{\mathcal{S}}P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x-\int_{\mathcal{S}}P_{h}^{\mathcal{E}}(x|s_{h},a_{h})\log\frac{P_{h}^{\mathcal{E}}(x|s_{h},a_{h})}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x\bigg{|}$
$+\bigg{|}\int_{[0,1]}R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x-\int_{[0,1]}R_{h}^{\mathcal{E}}(x|s_{h},a_{h})\log\frac{R_{h}^{\mathcal{E}}(x|s_{h},a_{h})}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|s_{h},a_{h})}\mathrm{d}x\bigg{|}.$ (A.30)
Let $o=(s_{h},a_{h})$. For the first term in Eqn. (A.4), we have the following bound
$\bigg{|}\int_{\mathcal{S}}P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\mathcal{E}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
$\leq\bigg{|}\int_{\mathcal{S}}P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
$\qquad+\bigg{|}\int_{\mathcal{S}}P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-P_{h}^{\mathcal{E}}(x|o)\log\frac{P_{h}^{\mathcal{E}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
@@
$\leq\int_{\mathcal{S}}\left|P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)-P_{h}^{\mathcal{E}}(x|o)\right|\cdot\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\bigg{|}\mathrm{d}x+\int_{\mathcal{S}}B\cdot\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
@@
$\leq B\cdot\left(1+\log\frac{B}{\beta}\right)\int_{\mathcal{S}}\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
$\leq\frac{(1+\log(B/\beta))\epsilon}{2H^{2}},$ (A.31)